OpenStudy (anonymous):

Solve the equation by completing the square: x2 + 8 = 6x Enter the smallest of the two zeros.

OpenStudy (zzr0ck3r):

we have x^2-6x+8 = 0 take half the middle term so -3 and move the 8 to the RHS and add half the middle term coeffecient squared to the RHS (x-3)^2 = -8+3^2 can you solve it from here?

OpenStudy (anonymous):

I though the 8 would be on the other side of the equal sign?

OpenStudy (markrijckenberg):

zzro0ck3r gave the right answer. This means there are not ONE, but TWO values/solutions for x.

OpenStudy (anonymous):

I know there are two values it's just confusing because I can't figure out how to do it cause apparently I'm doing it wrong

OpenStudy (zzr0ck3r):

its hard for me to explain how to complete the square on here. note: I took half the middle term (including its sign) and said (x+(half the middle term))^2 and now I want only this part on the LHS so I moved the 8 to the RHS. Then you need to add half the middle term squared to the RHS. (notice the sign does not matter once you square it) then solve for x

OpenStudy (zzr0ck3r):

when I say the middle term, I mean just the coefficient.

OpenStudy (zzr0ck3r):

aslo, you want no coefficients on the 1st term so devide them away first so if we have ax^2 + bx + c we get x^2+(b/a)x+(c/a) (note: a cant equal 0) then (x+(b/(2a)))^2 = -(c/a)+ (b/(2a))^2 then solve for x